# The Angular Momentum Penrose Inequality

### v2

@@ 4.3 Axisymmetric Setting
  • Verification of (P1): This is Han–Khuri [29, Proposition 4.5]. Specifically, for initial data $(M,g,K)$ satisfying DEC with a stable outermost MOTS $\Sigma$, the unperturbed Jang equation $\mathcal{J}_{0}[f]=0$ admits a solution $f_{0}$ with blow-up asymptotics $f_{0}(s,y)=C_{0}\ln s^{-1}+A_{0}(y)+O(s^{\alpha})$ where $C_{0}=|\theta^{-}|/2>0$ is determined by the inner null expansion $\theta^{-}=H_{\Sigma}-\mathrm{tr}_{\Sigma}K<0$. The exponent $\alpha>0$ depends on the spectral gap of the MOTS stability operator; for strictly stable MOTS, $\alpha=\min(1,2\sqrt{\lambda_{1}(L_{\Sigma})})$ where $\lambda_{1}(L_{\Sigma})>0$ is the principal eigenvalue.
  • Verification of (P2): This follows from Lockhart–McOwen [36, Theorem 7.4] combined with the Fredholm theory for asymptotically cylindrical operators developed by Melrose [40, Chapter 5]. We provide a detailed justification of the coercivity estimate.
@@
    Since $V\geq 0$ for stable MOTS (the MOTS stability condition states $\int_{\Sigma}|\nabla\psi|^{2}+(|A_{\Sigma}|^{2}+\mathrm{Ric}_{g}(\nu,\nu))\psi^{2}\geq 0$ for all test functions $\psi$, which implies $V\geq 0$ pointwise for stability with respect to all variations), we have:
@@
    Step (iv): Fredholm property. For $\beta$ in the valid range (not equal to any indicial root), [36, Theorem 1.1] implies $L_{0}:W^{2,2}_{\beta}\to W^{0,2}_{\beta}$ is Fredholm of index zero. The index is zero because the number of positive roots in $(0,\beta)$ equals the number of negative roots in $(\beta,0)$ (both are zero for $\beta\in(-\gamma_{0},0)$).
    Step (v): Kernel triviality. Suppose $L_{0}v=0$ with $v\in W^{2,2}_{\beta}$. Since $\beta<0$, we have $v\to 0$ as $t\to\infty$. An energy argument (multiply by $v$ and integrate) combined with the stability inequality shows $\int|\nabla v|^{2}+Vv^{2}\geq 0$. The boundary conditions and maximum principle force $v\equiv 0$. This kernel triviality is the key consequence of MOTS stability: elements of $\ker(L_{0})$ would correspond to infinitesimal deformations of the MOTS that preserve the marginally trapped condition, i.e., Jacobi fields. By [8, Proposition 3.2], stability of $\Sigma$ excludes non-trivial $L^{2}$-Jacobi fields.
@@
Lemma 4.17 (Indicial Roots for Twisted Jang Operator).
@@
where $\rho^{3}\omega=d\Omega$. By standard elliptic regularity [28, Theorem 8.32], $\Omega\in C^{2,\alpha}(\overline{\mathcal{Q}})$, which implies:
@@
Step 3: Indicial roots are unchanged. By [36, Theorem 6.1], the indicial roots of an elliptic operator $L$ on a manifold with cylindrical ends are determined by the translation-invariant limit operator $L_{\infty}$ obtained by taking $t\to\infty$. Since $D\mathcal{T}|_{f}=O(e^{-t})$ decays exponentially (with explicit rate from (36)), it does not contribute to $L_{\infty}$:
@@
(see [14, Section 3.2]). Thus $|\gamma_{1}|=\sqrt{\lambda_{1}}\geq\sqrt{8\pi/A}$.
Lockhart–McOwen condition. The theory in [36, Theorem 1.1] requires:
@@
Step 4: Fredholm property. By [36, Theorem 1.1], $L:W^{k,2}_{\beta}\to W^{k-2,2}_{\beta}$ is Fredholm if and only if $\beta$ is not an indicial root. The Fredholm index depends only on the indicial roots and their multiplicities. Since $L_{\mathrm{axi}}$ and $L_{0}$ have the same indicial roots, they have the same Fredholm index.
@@
Step 5b: Connection to MOTS stability operator. Following Andersson–Metzger [8, Section 3], we observe that the limiting potential $\mathcal{V}$ is related to the MOTS stability operator (Definition 4.5).
@@
Step 5d: Using MOTS stability. The MOTS stability condition $\lambda_{1}(L_{\Sigma})\geq 0$ means:
@@
for all $\psi\in C^{\infty}(\Sigma)$. Equivalently, $\int_{\Sigma}\mathcal{V}_{\Sigma}\psi^{2}\leq\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}$ where $\mathcal{V}_{\Sigma}=|A_{\Sigma}|^{2}+\mathrm{Ric}(\nu,\nu)\geq 0$ by stability.
@@
independent of $\kappa,\delta$. Interior Schauder estimates (using DEC to prevent interior gradient blow-up) give $C^{2,\alpha}_{\text{loc}}$ compactness. Taking a diagonal subsequence as $\kappa\to 0,\delta\to 0$:
$f_{\kappa,\delta}\to f\quad\text{in }C^{2,\alpha}_{\text{loc}}(M\setminus\Sigma),$
@@
  • The $\rho^{2}$ scaling of the twist term: $\mathcal{T}\propto\rho^{2}$, which vanishes at the poles where $\rho=0$ (Lemmas 4.7 and 4.9).
@@ 5.1 The Conformal Equation
    Asymptotically flat end: Weighted Hölder spaces $C^{k,\alpha}_{-\tau}$ with polynomial weight $r^{-\tau}$ (Definition 4.2);
@@
These frameworks are compatible on the transition region $\{R_{0}\leq r\leq 2R_{0}\}$ (equivalently $\{0\leq t\leq T_{0}\}$) in the following sense: by Sobolev embedding, $W^{k+1,2}_{\beta}\hookrightarrow C^{k,\alpha}$ locally, and both norms are equivalent (up to constants depending on $R_{0}$) on the compact overlap region. This allows elliptic estimates to be “glued” across the transition using standard partition-of-unity arguments. The key point is that the Fredholm index is determined by the asymptotic behavior at both ends, not the transition region.
@@
where $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}_{\bar{g}}\geq 0$ is the transverse-traceless contribution (encoding rotation). The negative sign in front of $\Lambda_{J}$ ensures that the conformal scalar curvature $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$.
@@
    Positivity: Since $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}\geq 0$ and $\phi>0$, we have $R_{\tilde{g}}\geq 0$ automatically.
@@
    Strict positivity: $R_{\tilde{g}}>0$ where $\sigma^{TT}\neq 0$, i.e., where the data has non-trivial gravitational radiation.
@@
The convention matches the standard Lichnerowicz equation $8\Delta\phi+R\phi=0$ (for $R_{\tilde{g}}=0$), with the $\Lambda_{J}\phi^{-7}$ term producing positive conformal scalar curvature.
Key Formula: Conformal Scalar Curvature Derivation The conformal scalar curvature $R_{\tilde{g}}$ is determined algebraically from the AM-Lichnerowicz equation: 1. Standard conformal transformation ($\tilde{g}=\phi^{4}\bar{g}$ in dimension 3): $R_{\tilde{g}}=\phi^{-5}\bigl(\phi\cdot R_{\bar{g}}-8\Delta_{\bar{g}}\phi\bigr)$ 2. AM-Lichnerowicz equation (Definition 5.3): $8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=\Lambda_{J}\phi^{-7}$ 3. Substitution: Replace $(\phi R_{\bar{g}}-8\Delta_{\bar{g}}\phi)$ by $\Lambda_{J}\phi^{-7}$: $R_{\tilde{g}}=\phi^{-5}\cdot\Lambda_{J}\phi^{-7}=\Lambda_{J}\phi^{-12}$ 4. Conclusion: Since $\Lambda_{J}=\frac{1}{8}|\sigma^{TT}|^{2}\geq 0$ and $\phi>0$, we have: $\boxed{R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0}$ This is independent of whether $\phi\leq 1$ or $\phi>1$. The bound $R_{\tilde{g}}\geq 0$ is the only curvature condition needed for AMO monotonicity.
Lemma 5.5 (Fredholm Property).
The linearization $L:=-8\Delta_{\bar{g}}+R_{\bar{g}}$ of the operator in (40) at $\phi=1$ is Fredholm
$L:W^{2,2}_{\beta}(\bar{M})\to L^{2}_{\beta}(\bar{M})$
of index zero for $\beta\in(-1,0)$ not equal to any indicial root.

### v3

@@ 4.3 Axisymmetric Setting
  • Verification of (P1): This is Han–Khuri [29, Proposition 4.5]. Specifically, for initial data $(M,g,K)$ satisfying DEC with a stable outermost MOTS $\Sigma$, the unperturbed Jang equation $\mathcal{J}_{0}[f]=0$ admits a solution $f_{0}$ with blow-up asymptotics $f_{0}(s,y)=C_{0}\ln s^{-1}+\mathcal{A}_{0}(y)+O(s^{\alpha})$ where $C_{0}=|\theta^{-}|/2>0$ is determined by the inner null expansion $\theta^{-}=H_{\Sigma}-\mathrm{tr}_{\Sigma}K<0$. The exponent $\alpha>0$ depends on the spectral gap of the MOTS stability operator; for strictly stable MOTS, $\alpha=\min(1,2\sqrt{\lambda_{1}(L_{\Sigma})})$ where $\lambda_{1}(L_{\Sigma})>0$ is the principal eigenvalue.
  • Verification of (P2): This follows from Lockhart–McOwen [35, Theorem 7.4] combined with the Fredholm theory for asymptotically cylindrical operators developed by Melrose [41, Chapter 5]. We provide a detailed justification of the coercivity estimate.
@@
    The MOTS stability inequality implies the quadratic form associated to $\Delta_{\Sigma}-V$ is nonnegative on $H^{1}(\Sigma)$, i.e.
    $\int_{\Sigma}|\nabla\psi|^{2}+V\,\psi^{2}\,dA\geq 0\quad\text{for all }\psi\in C^{\infty}(\Sigma).$
    We will only use this inequality (not any pointwise sign for $V$):
@@
    Step (iv): Fredholm property. For $\beta$ in the valid range (not equal to any indicial root), [35, Theorem 1.1] implies $L_{0}:W^{2,2}_{\beta}\to W^{0,2}_{\beta}$ is Fredholm of index zero. The index is zero because the number of positive roots in $(0,\beta)$ equals the number of negative roots in $(\beta,0)$ (both are zero for $\beta\in(-\gamma_{0},0)$).
    Step (v): Kernel triviality. Suppose $L_{0}v=0$ with $v\in W^{2,2}_{\beta}$. Since $\beta<0$, we have $v\to 0$ as $t\to\infty$. An energy argument (multiply by $v$ and integrate) combined with the stability inequality shows $\int|\nabla v|^{2}+Vv^{2}\geq 0$. The boundary conditions and maximum principle force $v\equiv 0$. This kernel triviality is the key consequence of MOTS stability: elements of $\ker(L_{0})$ would correspond to infinitesimal deformations of the MOTS that preserve the marginally trapped condition, i.e., Jacobi fields. By [7, Proposition 3.2], stability of $\Sigma$ excludes non-trivial $L^{2}$-Jacobi fields.
@@
Lemma 4.18 (Fredholm Theory for Twisted Jang Operator).
@@
where $\rho^{3}\omega=d\Omega$. By standard elliptic regularity [28, Theorem 8.32], $\Omega\in C^{2,\beta}(\overline{\mathcal{Q}})$, which implies:
@@
Step 3: Indicial roots are unchanged. By [35, Theorem 6.1], the indicial roots of an elliptic operator $L$ on a manifold with cylindrical ends are determined by the translation-invariant limit operator $L_{\infty}$ obtained by taking $t\to\infty$. Since $D\mathcal{T}|_{f}=O(e^{-t})$ decays exponentially (with explicit rate from (36)), it does not contribute to $L_{\infty}$:
@@
(see [13, Section 3.2]). Thus $|\gamma_{1}|=\sqrt{\lambda_{1}}\geq\sqrt{8\pi/A}$.
Lockhart–McOwen condition. The theory in [35, Theorem 1.1] requires:
@@
Step 4: Fredholm property. By [35, Theorem 1.1], $L:W^{k,2}_{\beta}\to W^{k-2,2}_{\beta}$ is Fredholm if and only if $\beta$ is not an indicial root. The Fredholm index depends only on the indicial roots and their multiplicities. Since $L_{\mathrm{axi}}$ and $L_{0}$ have the same indicial roots, they have the same Fredholm index.
@@
Step 5b: Connection to MOTS stability operator. Following Andersson–Metzger [7, Section 3], we observe that the limiting potential $\mathcal{V}$ is related to the MOTS stability operator (Definition 4.4).
@@
Step 5d: Using MOTS stability. The MOTS stability condition $\lambda_{1}(L_{\Sigma})\geq 0$ yields the (quadratic-form) stability inequality
@@
for all $\psi\in C^{\infty}(\Sigma)$. Equivalently, writing $\mathcal{V}_{\Sigma}:=|A_{\Sigma}|^{2}+\mathrm{Ric}_{g}(\nu,\nu)$, we have
$\int_{\Sigma}\mathcal{V}_{\Sigma}\,\psi^{2}\,d\sigma\leq\int_{\Sigma}|\nabla_{\Sigma}\psi|^{2}\,d\sigma\quad\text{for all }\psi\in C^{\infty}(\Sigma),$
and we will use only this integral inequality (not any pointwise sign for $\mathcal{V}_{\Sigma}$).
@@
independent of $\kappa,\delta$. Interior Schauder estimates (using DEC to prevent interior gradient blow-up) give $C^{2,\beta}_{\text{loc}}$ compactness. Taking a diagonal subsequence as $\kappa\to 0,\delta\to 0$:
$f_{\kappa,\delta}\to f\quad\text{in }C^{2,\beta}_{\text{loc}}(M\setminus\Sigma),$
@@
  • The $\rho^{2}$ scaling of the twist term: $\mathcal{T}\propto\rho^{2}$, which vanishes at the poles where $\rho=0$ (Lemmas 4.6 and 4.8).
@@ 5.1 The Conformal Equation
    Asymptotically flat end: Weighted Hölder spaces $C^{k,\beta}_{-\tau}$ with polynomial weight $r^{-\tau}$ (Definition 4.1);
@@
These frameworks are compatible on the transition region $\{R_{0}\leq r\leq 2R_{0}\}$ (equivalently $\{0\leq t\leq T_{0}\}$) in the following sense: by Sobolev embedding, $W^{k+1,2}_{\beta}\hookrightarrow C^{k,\beta}$ locally, and both norms are equivalent (up to constants depending on $R_{0}$) on the compact overlap region. This allows elliptic estimates to be “glued” across the transition using standard partition-of-unity arguments. The key point is that the Fredholm index is determined by the asymptotic behavior at both ends, not the transition region.
@@
where $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}\geq 0$ is the Kerr deviation contribution (Definition 1.11). The negative sign in front of $\Lambda_{J}$ ensures that the conformal scalar curvature $R_{\tilde{g}}=\Lambda_{J}\phi^{-12}\geq 0$.
Key property: For Kerr initial data, $\Lambda_{J}=0$ (since $\mathcal{S}_{(g,K)}=0$), and the equation reduces to the standard Lichnerowicz equation $8\Delta_{\bar{g}}\phi+R_{\bar{g}}\phi=0$.
@@
    Positivity: Since $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}\geq 0$ and $\phi>0$, we have $R_{\tilde{g}}\geq 0$ automatically.
@@
    Strict positivity: $R_{\tilde{g}}>0$ where $\mathcal{S}_{(g,K)}\neq 0$, i.e., where the data deviates from Kerr geometry.
  5. (v)
    Equality case: For Kerr data, $\Lambda_{J}=0$, so $R_{\tilde{g}}=0$ and the monotonicity integrand vanishes.
@@
The convention matches the standard Lichnerowicz equation $8\Delta\phi+R\phi=0$ (for $R_{\tilde{g}}=0$), with the $\Lambda_{J}\phi^{-7}$ term producing positive conformal scalar curvature for non-Kerr data.
Lemma 5.5 (Well-Definedness of $\Lambda_{J}$).
The angular momentum source term $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}$ is a well-defined, coordinate-independent, non-negative scalar function on any asymptotically flat, axisymmetric vacuum initial data $(M^{3},g,K)$. We provide two equivalent constructions: an intrinsic algebraic definition (simpler, self-contained) and a PDE-based extension (connecting to Mars–Simon theory).
Construction A: Intrinsic Algebraic Definition (Primary).
  1. (i)
    Electric and magnetic Weyl tensors: Define algebraically from $(g,K)$:
    $E_{ij}$ $:=R_{ij}-\tfrac{1}{3}Rg_{ij}+(\mathrm{tr}K)K_{ij}-K_{ik}K^{k}{}_{j}$,
    $B_{ij}$ $:=\epsilon_{i}{}^{kl}\nabla_{k}K_{lj}.$
    These are intrinsic to (g,K)(g,K)—no embedding or evolution required.
  2. (ii)
    Reference Kerr tensors via asymptotic expansion: For asymptotically flat data with ADM mass $M$ and Komar angular momentum $J$, define the reference Kerr Weyl tensors $E^{\mathrm{Kerr}}_{ij}(M,J)$ and $B^{\mathrm{Kerr}}_{ij}(M,J)$ by the explicit asymptotic series:
    $E^{\mathrm{Kerr}}_{ij}$ $=\frac{M}{r^{3}}\left(\delta_{ij}-3\hat{r}_{i}\hat{r}_{j}\right)+\frac{3Ma^{2}}{r^{5}}\left(\text{spin-2 harmonics}\right)+O(r^{-6}),$
    $B^{\mathrm{Kerr}}_{ij}$ $=\frac{3Ma}{r^{4}}\epsilon_{(i}{}^{kl}\hat{r}_{j)}\hat{r}_{k}\hat{z}_{l}+O(r^{-5}),$
    where $a=J/M$, $\hat{r}=x/r$, and $\hat{z}$ is the symmetry axis unit vector. Within the class of smooth, symmetric trace-free tensors admitting the stated Kerr asymptotic expansion (and with the prescribed leading-order coefficients), this fixes the reference tensors uniquely.
  3. (iii)
    Kerr deviation tensor: Define pointwise:
    $\mathcal{S}_{(g,K),ij}:=(E_{ij}-E^{\mathrm{Kerr}}_{ij})+i(B_{ij}-B^{\mathrm{Kerr}}_{ij}).$
    This is well-defined for $r>r_{0}$ (exterior region) where the asymptotic expansions converge.
  4. (iv)
    Angular momentum source: Define
    $\Lambda_{J}:=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}=\frac{1}{8}\left(|E-E^{\mathrm{Kerr}}|^{2}+|B-B^{\mathrm{Kerr}}|^{2}\right).$
Key properties (immediate from the construction):
  1. (a)
    $\Lambda_{J}\geq 0$ everywhere (squared norm);
  2. (b)
    $\Lambda_{J}$ is coordinate-independent (tensor norm);
  3. (c)
    $\Lambda_{J}=O(r^{-4-2\tau})$ for asymptotically flat data with decay $\tau>1/2$;
  4. (d)
    For Kerr slices: $E=E^{\mathrm{Kerr}}$ and $B=B^{\mathrm{Kerr}}$ exactly, so $\Lambda_{J}=0$;
  5. (e)
    $\Lambda_{J}=0$ iff $(M,g,K)$ is a Kerr slice (Theorem G.14).
Construction B: PDE Extension (for Interior Region). For the strong-field region (e.g., near the MOTS where the asymptotic expansion may not converge), we extend $E^{\mathrm{Kerr}}$ and $B^{\mathrm{Kerr}}$ to all of $M$ via:
  1. (i)
    Constraint propagation: The constraint equations for vacuum data imply the Codazzi–Mainardi identity:
    $\nabla^{j}E_{ij}=\epsilon_{ijk}K^{jl}B^{k}{}_{l},\qquad\nabla^{j}B_{ij}=-\epsilon_{ijk}K^{jl}E^{k}{}_{l}$.
    This is a determined system (not elliptic in the standard sense, but constrained by the Bianchi identity).
  2. (ii)
    Well-posedness via harmonic analysis: Following [86, 87], one can construct reference tensors $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ satisfying the Codazzi–Mainardi system and the prescribed Kerr asymptotics. Moreover, within the corresponding asymptotic/regularity class (as in [86, 87]), this extension is unique. The proof uses:
    • Decomposition into spherical harmonics on large spheres;
    • The Codazzi–Mainardi system determines the radial evolution of each harmonic mode;
    • Uniqueness follows from the decay conditions at infinity (i.e., the only decaying solution of the homogeneous system in that class is the trivial one).
  3. (iii)
    Regularity: The extended tensors $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ are smooth on $M\setminus\Sigma$, and extend (at least) continuously up to $\Sigma$ under the regularity assumptions imposed on $(g,K)$; we only use such regularity as needed to make $\Lambda_{J}$ well-defined and integrable.
Remark 5.6 (Rigorous Well-Posedness of the Interior Extension).
We provide a complete proof that the reference tensors $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ extend uniquely to the interior region. The argument has three parts.
Part 1: System structure. The Codazzi–Mainardi system for symmetric trace-free tensors $(E,B)$ is:
$\nabla^{j}E_{ij}$ $=\epsilon_{ijk}K^{jl}B^{k}{}_{l}=:F_{i}(E,B,K),$ (41)
$\nabla^{j}B_{ij}$ $=-\epsilon_{ijk}K^{jl}E^{k}{}_{l}=:G_{i}(E,B,K).$ (42)
Writing $(E,B)=(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$, the system becomes a linear first-order system with coefficients depending on $(g,K)$.
Part 2: Spherical harmonic decomposition. On a sphere $S_{r}$ of radius $r$, decompose:
$E_{ij}|_{S_{r}}=\sum_{\ell\geq 2}\sum_{|m|\leq\ell}E_{\ell m}(r)Y^{\ell m}_{ij}(\theta,\phi),\quad B_{ij}|_{S_{r}}=\sum_{\ell\geq 2}\sum_{|m|\leq\ell}B_{\ell m}(r)Y^{\ell m}_{ij}(\theta,\phi),$
where $Y^{\ell m}_{ij}$ are the symmetric trace-free tensor spherical harmonics. The Codazzi–Mainardi system (41)–(42) becomes a system of ODEs for the radial coefficients:
$\frac{d}{dr}\begin{pmatrix}E_{\ell m}\\ B_{\ell m}\end{pmatrix}=A_{\ell}(r)\begin{pmatrix}E_{\ell m}\\ B_{\ell m}\end{pmatrix}+\text{(lower order in }\ell),$ (43)
where $A_{\ell}(r)$ is a matrix depending on $(\ell,r,g,K)$.
Part 3: Well-posedness via ODE theory.
  1. (a)
    Boundary data at infinity: The Kerr asymptotics determine $(E^{\mathrm{Kerr}}_{\ell m},B^{\mathrm{Kerr}}_{\ell m})|_{r=\infty}$ for each mode. Explicitly:
    $E^{\mathrm{Kerr}}_{2,0}(r)=\frac{M}{r^{3}}(1+O(r^{-2})),\quad B^{\mathrm{Kerr}}_{2,0}(r)=\frac{3Ma}{r^{4}}(1+O(r^{-1})),$
    with higher modes decaying faster.
  2. (b)
    Inward integration: Given boundary values at $r=r_{0}$ (sufficiently large), the ODE system (43) has a unique solution by Picard–Lindelöf on any interval on which the coefficients remain regular. In particular, it extends uniquely for $r>0$ away from the axis (and away from any other region where the chosen radial foliation may degenerate).
  3. (c)
    Axis regularity: For axisymmetric data ($m=0$ modes only), the tensor spherical harmonics have the form $Y^{\ell 0}_{ij}\propto P_{\ell}(\cos\theta)\times(\text{angular structure})$. The Legendre functions $P_{\ell}$ are smooth at the poles $\theta=0,\pi$, so $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ extend smoothly to the axis.
  4. (d)
    Near-MOTS behavior: As $r\to r_{\mathrm{MOTS}}$, the coefficients $A_{\ell}(r)$ remain bounded (they depend on $g,K$, which are smooth up to the MOTS). The ODE solution $(E^{\mathrm{Kerr}}_{\ell m},B^{\mathrm{Kerr}}_{\ell m})$ therefore extends continuously to $r=r_{\mathrm{MOTS}}$.
[…]
Explicit interior construction procedure: The reference Kerr tensors $E^{\mathrm{Kerr}}_{ij}$ and $B^{\mathrm{Kerr}}_{ij}$ can be extended to the interior region $\{r<r_{0}\}$ via the following steps:
  1. (a)
    Boundary data: On a large sphere $S_{r_{0}}$, compute the asymptotic values $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})|_{S_{r_{0}}}$ from the explicit Kerr formulas.
  2. (b)
    Inward integration: Solve the Codazzi–Mainardi system inward from $S_{r_{0}}$ toward the MOTS. Since the system is first-order in the radial direction (after harmonic decomposition), this is a well-posed ODE for each harmonic mode.
  3. (c)
    Regularity at axis: The axisymmetry condition $\mathcal{L}_{\eta}E^{\mathrm{Kerr}}=\mathcal{L}_{\eta}B^{\mathrm{Kerr}}=0$ constrains the harmonic modes to those compatible with axis regularity (only $m=0$ azimuthal modes for the scalar quantities).
The resulting $(E^{\mathrm{Kerr}},B^{\mathrm{Kerr}})$ are smooth throughout $M\setminus\Sigma$ and satisfy the Codazzi–Mainardi equations by construction.
extbfEquivalence: Constructions A and B agree in the overlap region $\{r>r_{0}\}$ because both satisfy the same Codazzi–Mainardi system and the same Kerr asymptotics (hence coincide by uniqueness in the relevant asymptotic class). The extension then supplies values in the interior.